I have 2 questions. One I need to see if I got right and the other I have no clue about. Thanks to anyone willing to assist.
Solve the following equation for x.
23x^2 + 17x= -11
23x^2 + 17x + 11=0
I get that there is no solution b/c this can't be factored. Correct?
The other is 1 + radical(x-5) =0
I dont have the radical symbol. Just to be clear, x-5 is under the radical.

Found 2 solutions by stanbon, uma:
Answer by stanbon(48535)

Solve the following equation for x.
23x^2 + 17x= -11
23x^2 + 17x + 11=0
I get that there is no solution b/c this can't be factored. Correct?
The determinant, b^2-4ac, is negative; that tells you the factors
involve complex numbers. It can be factored but not in the Real
Number System.
-------------------
The other is 1 + radical(x-5) =0
I dont have the radical symbol. Just to be clear, x-5 is under the radical.
sqrt(x-5)=-1
The square root cannot be negative; so the equation has no solution
in the Real Number System.
Cheers,
Stan H.

Answer by uma(370)

Yes you are right.
23x^2 + 17x + 11 = 0 cannot be factored.
Even if you try to use the formula , the solution is not real.
The second one is 1 + Sqrt(x-5) = 0
==> sqrt (x-5) = -1 [ adding -1 to both the sides]
==> x - 5 = 1 [Squaring both the sides]
==> x = 6 [Adding 5 to both the sides]
So x = 6 and the given eqn is satisfied only for negative sqrt.
Regards,
Uma